1. The supply and demand for saffron (a spice) are described by the following equations: Supply: ??=6?−12 Demand: ??=252−2? Now consider the policies proposed by each candidate:

c) Candidate A, who hates saffron, proposes banning the spice altogether. Find the CS, PS, and TS. Who would bear the larger burden of this policy – the buyer or sellers of saffron?

d) Candidate B likes saffron but feels it is too expenses. Thus B’s policy proposal is to place a price ceiling of $27 on saffron. How many ounces of saffron would be sold? Illustrate and calculate the impact of this policy on the CS, PS, and TS. Would consumers benefit from the price ceiling? Support your answer with no more than two sentences.

e) Candidate C prefers to set a maximum number of ounces of saffron that can be exchanged in the market at 150 ounces. Draw a graph illustrating the market supply curve that would reflect the impact of this policy. How many ounces of saffron would be sold? Find the CS, PS, and TS under this policy.

f) Which policy yields the highest level of welfare for society? Is the best of these policies preferable to the free market equilibrium? Support your answers using the amount of welfare loss from each policy.
Answer #1

26 6x33-12 19.8. 12 12.6 3 6 25 2 Ps 2 2 83.3 86282 ed w 6.127)-1 S Seanned = 15D CamScanner25 224 =138 126 51 32 2.3 Prile Cejuus2 2 15D 6 98 252 At &2150, &d- 25 2-2p 156 = 252-2p→2?=252-150 cs- LI)26-5))(15り (5)-2712-6 gS 23 21 pD 2 156 252 8 852-P15o 0 L-X-7-5- 6 25 25,0 51-2) 15o 4구5 Sca CamScanner432 貝 .umda.. Pa madat-equalǐとmum ts3.L.532 CS Scanned with CamScanner